OpenStudy (thedj4jc):

What are the excluded values? (problem in the replies) x = 0 and 3 x = 0 and 2 x = -3 and 3 none of the above

OpenStudy (thedj4jc):

\[\frac{ 2 }{ x(x-3) }\]

OpenStudy (thedj4jc):

2. What are the excluded values? \[\frac{ y+5 }{ y ^{2}+4y-32 }\] y = -5 and 4 y = 5 and 32 y = 4 and -8 none of the above

OpenStudy (anonymous):

Well for the first one, it's basically asking "which values will make the bottom equal 0?" At that point, you can just plug in the values to see which ones make it equal 0.

OpenStudy (thedj4jc):

That'd be A, cause 3-3=0 and 0*0=0

OpenStudy (anonymous):

Exactly, so A would be your answer. The same concept applies to the second one. Just plug in the numbers to see which ones equal 0.

OpenStudy (thedj4jc):

Ok, thank you!!!!

OpenStudy (anonymous):

No problem :)